The pH of a sample of human blood was measured to be 7.41 at 25 oC. What is the hydroxide ion [OH-(aq)] concentration?
1
Expert's answer
2019-08-21T05:31:33-0400

pH =7.41

pOH = 14- pH = 6.59

[OH-(aq)]  = 10-pOH= 2.57*10-7 mol/dm3